1. The raw scores of 20 students of Whitesands School who took part in an examination in Economics are given below. The pass marks is
40%
38 28 70 43
39 20 64 66
12 46 52 53
20 34 48 69
18 20 64 34
(a) What is the mean score of the students’ marks?
(b) How many students passed the examination?
© What percentage of the students failed the examination?
(d) What is the range of the scores?
(e) How many students scored below the mean score?
1
Expert's answer
2019-04-30T09:21:38-0400

(a) The mean score of the students’ marks is:

X = (38 + 28 + 70 + 43 + 39 + 20 + 64 + 66 + 12 + 46 + 52 + 53 + 20 + 34 + 48 + 69 + 18 + 20 + 64 + 34) /20 = 41.9%.

(b) 10 of 20 students passed the examination, because their score is higher than 40%.

(c) The percentage of the students that failed the examination is: (20 - 10)/20×100% = 50%.

(d) What is the range of the scores is 70 - 12 = 58.

(e) 10 students scored below the mean score.